# IDDM: Bridging Synthetic-to-Real Domain Gap from Physics-Guided Diffusion for Real-world Image Dehazing

### v1

@@ Abstract
Due to the domain gap between real-world and synthetic hazy images, current data-driven dehazing algorithms trained on synthetic datasets perform well on synthetic data but struggle to generalize to real-world scenarios. To address this challenge, we propose Image Dehazing Diffusion Models (IDDM), a novel diffusion process that incorporates the atmospheric scattering model into noise diffusion. IDDM aims to use the gradual haze formation process to help the denoising Unet robustly learn the distribution of clear images from the conditional input hazy images. We design a specialized training strategy centered around IDDM. Diffusion models are leveraged to bridge the domain gap from synthetic to real-world, while the atmospheric scattering model provides physical guidance for haze formation. During the forward process, IDDM simultaneously introduces haze and noise into clear images, and then robustly separates them during the sampling process. By training with physics-guided information, IDDM shows the ability of domain generalization, and effectively restores the real-world hazy images despite being trained on synthetic datasets. Extensive experiments demonstrate the effectiveness of our method through both quantitative and qualitative comparisons with state-of-the-art approaches.
@@ 1 Introduction
Figure 1: Haze formation and the diffusion forward process share a common characteristic: both degrade the clarity of a clear image, whether by introducing noise or haze. This paper addresses real-world image dehazing by integrating the image formation process into the diffusion process.
Due to the attenuation of light in the atmosphere, images captured in hazy conditions often suffer from blurriness. These degradations not only affect visual quality but also significantly impair the performance of computer vision systems in critical applications such as autonomous driving and surveillance. To address the obvious harm to machine vision systems, extensive research [1, 2, 31, 30, 32, 14] has been conducted in recent years on image dehazing.
Prior-based methods [14, 1] frequently lack robustness in scenarios where prior information is unavailable. In contrast, data-driven deep learning methods [56, 2] have proven to be more effective in handling complex environmental conditions. Current deep learning approaches for image dehazing [53, 7] are often trained on synthetic datasets [21, 24] with paired hazy and haze-free images, as acquiring real-world paired hazy data is challenging. However, models trained exclusively on synthetic data often exhibit poor generalization capabilities when applied to real-world images, especially under complex and variable weather conditions. This domain gap significantly limits the practical utility of these algorithms in real-world applications.
@@
To implement this progressive learning approach, we turn to diffusion models, which have demonstrated remarkable success in various image generation and restoration tasks. We find that diffusion models offer an ideal foundation for this setting for several compelling reasons. The denoising diffusion models, which also contain inherent progressive learning strategies through their step-by-step noise addition and removal processes, excel in data distribution modeling, which can address the cross-domain challenge between synthetic and real images by learning the natural distribution of clear images. Moreover, there exists a remarkable parallel between the forward process of DDPM [15] and the physical process of haze formation in real-world scenarios.
As illustrated in Fig. 1, in a forward diffusion process, a clear image gradually becomes pure noise over time, while in the haze formation process, it gradually accumulates haze as depth increases. Both processes share a similar trajectory: the image progressively deteriorates, either into noisy or hazy. This natural correspondence provides a theoretical basis for integrating physical models into generative frameworks and makes diffusion models particularly suitable as a foundation for image dehazing.
@@
To overcome this limitation, we utilize their similarities to deduce a novel diffusion model (IDDM) specifically for image dehazing, which explicitly models the formation of haze through the diffusion process. We model haze as a timestep-dependent haze component $\mathbf{h}_{t}$ to integrate physical principles into the diffusion process. A unique bi-directional enhancement framework is created around it, enabling denoising and dehazing processes to mutually promote each other. $\mathbf{h}_{t}$ forms the foundation of our novel image dehazing diffusion model and serves as a conditional vector guiding the diffusion model’s training, enabling accurate noise prediction across varying haze densities. Simultaneously, intermediate images with varying levels of noise and haze from the diffusion process are utilized to train the model to estimate haze under diverse conditions at each timestep.
Our method achieves substantial performance improvements on real-world datasets, demonstrating its effectiveness in bridging the synthetic-to-real domain gap. On the RTTS dataset [21], we observe improvements of 19.44% in Brisque, 10.68% in NIQE, 5.82% in PIQE, and 4.79% in NIMA metrics compared to previous methods. Furthermore, on the BEDDE dataset [54], our approach demonstrates significant improvements of 3.90% in VI, 0.8% in RI, and 2.97% in VSI metrics. These results validate the effectiveness of leveraging physical principles for real-world image dehazing.
@@
(1) We introduce Image Dehazing Diffusion Models (IDDM), a novel framework that integrates the atmospheric scattering model into diffusion processes. This integration allows IDDM to simultaneously leverage the distribution learning capabilities of diffusion models and the physical constraints of atmospheric scattering, creating a physically-informed approach to dehazing
(2) We introduce an IDDM training procedure that adapts to varying degrees of haze. Instead of solely focusing on learning the mapping from hazy to clear images, our model learns hazy formation at each time $t$ and the distribution of clear images.
@@ 2.1 Single Image Dehazing
Most prior-based image dehazing works originate from the atmospheric scattering model (ASM) [30, 19, 31]. He et al. [14] proposed the classical dark channel prior, which identifies that there exists at least one channel close to zero for each pixel. Berman et al. [1] introduced the haze-lines prior, which assumes that colors of a haze-free image can be approximated by a few hundred distinct colors. While these traditional prior-based methods [1, 14, 22, 10, 9, 44, 58, 16, 45] are efficient, they are not robust in dealing with complex scenes. Early deep learning methods adopt CNN-based networks [2, 37, 51] to calculate transmission map and airlight for single image dehazing.
Furthermore, networks that directly restore clear images from hazy images [8, 2, 20, 23, 34, 36] are designed without considering physical principles, which may lead to artifacts, color casts, poor domain generalization, and inability to address dense haze. To achieve a good performance for hazy images, another line of works [55, 47] incorporates ASM to help the networks learn the physics knowledge. Zheng et al. [55] reformulate ASM to learn the distribution of haze, with contrastive regularization adopted to solve the strong hazy images. These findings motivate us to investigate diffusion models as a powerful framework, leveraging their natural capacity for physical process modeling to achieve robust cross-domain image dehazing.
@@ 2.2 Diffusion Models
Diffusion-Based Methods [40] have garnered significant interest in the field of generation and restoration. The sampling process is conducted by a time-conditional additive noise estimator [15, 18] or a noise-conditional score function [42], both parameterized by deep neural networks. These models are then employed for stepwise denoising of samples from a noise distribution, ultimately obtaining samples from the data distribution.
Diffusion models have demonstrated exceptional performance in various image processing tasks, including image restoration [48, 12, 60, 12], image super-resolution [11], image deblurring [46], image editing, transfer, and generation [38, 39, 5, 25, 49, 52, 59], as well as 3D reconstruction [57, 17]. Additionally, conditional generative modeling from a score-based perspective has also been explored [26, 6]. These diffusion methods have played a crucial role in generative modeling, with continuous advancements and robust performance across a wide range of image processing tasks. These promising properities motivate us to leverage the diffusion model bridging the gap between synthetic and real-world domains, while also learning the physical process of haze formation with the guidance of the atmospheric scattering model.
@@ 3 Image Dehazing Diffusion Models
Given a data point sampled from a real data distribution $\mathbf{x}_{0}$, the forward diffusion process of DDPM at step $t$ can be explained as:
@@
where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\alpha_{i}$ denotes noise weight. In this equation, $\mathbf{x}_{t}$ represents the data at time step $t$, $\alpha_{t}$ is the scaling factor that controls noise, and $\boldsymbol{\epsilon}_{t}$ is Gaussian noise sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. This process closely resembles the generation of haze. A hazy image can be represented by the atmospheric scattering model (ASM) [30, 31] as:
$I=JT_{r}+A(1-T_{r}),$ (2)
where
$T_{r}=e^{-\beta z}.$ (3)
Here, $J$ denotes a clear image, $I$ denotes a hazy image, $A$ denotes the airlight, $z$ denotes depth, and $T_{r}$ denotes the transmission map. As denoted in Eq. (2) and Eq. (3), as the distance to the object plane increases, the haze part $A(1-T_{r})$ becomes denser until it reaches the camera plane. Inspired by their similarities that both processes progressively degrade through mathematically structured transformations, we integrate these two processes by modeling haze as a timestep-dependent component within the diffusion framework, and formulate image dehazing as a reverse diffusion process.
3.1 Forward Process
As shown in Fig. 2, the generative modeling framework of IDDM is based on the Denoising Diffusion Probabilistic Models [15]. Inspired by the idea of gradually adding noise to clear images, we simultaneously add noise and haze to the target distribution. Let a point on an image plane has a depth denoted as $Z$ in 3D world, with its reflectance intensity denoted as $\mathbf{x}_{0}$. The intensity of haze reaching the camera from this point can be as an integral, which is related to depth:
[…]
We take this as a diffusion process $\mathbf{h}_{t}$ ranging from depth 00 to depth $Z$, occurring over $T$ steps as:
[…]
Figure 2: The flowchart of IDDM process. Here, physics generation denotes Eq. (6), forward process denotes Sec. 3.1, and sampling process denotes Sec. 3.2.
We define the haze intensity at time $t$ as the integral of depth from zero to $\frac{t}{T}\cdot Z$:
[…]
Here, the variation of haze at different depths is represented by the time step $t$ in the diffusion process, which allows the gradual change of fog to be integrated into the diffusion model. $\mathbf{h}_{t}$ evolves synchronously with $\boldsymbol{\epsilon}_{t}$: as the noise $\boldsymbol{\epsilon}_{t}$ intensifies over time, the haze $\mathbf{h}_{t}$ also thickens progressively.
The change in haze between times $t-1$ and $t$ can be expressed as:
[…]
And $\mathbf{h}_{T}$ is the cumulative sum of $\Delta\mathbf{h}_{t}$ from $t=1$ to $T$:
$\mathbf{h}_{T}=\sum_{t=1}^{T}\Delta\mathbf{h}_{t},$ (8)
Therefore, the haze intensity in an image, $\mathbf{h}_{T}$, can be obtained by adding $\Delta\mathbf{h}_{t}$ over $T$ steps. To tailor the diffusion model for more effective image dehazing, we introduce a function $f$ to describe the haze intensity at each time step $t$. Based on the analysis of DDPM, we incorporate a haze formation term $f(\Delta\mathbf{h}_{t})$ into the model. This function $f$ should be related to $\Delta\mathbf{h}_{t}$ in the forward process from $t-1$ to $t$, and related to $\mathbf{h}_{t}$ in the forward process from 0 to $t$. We expect $f$ to have a compact mathematical form that exhibits the following properties:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x_{t-1}}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{t-1}+f(\Delta\mathbf{h}_{t}),$ (9)
which is a diffusion model that related to $\Delta\mathbf{h}_{t}$. Based on DDPM [15], $\mathbf{x}_{t}$ can not only be inferred from $\mathbf{x}_{t-1}$, but also can be expressed like that in Eq. (1), where $\mathbf{x}_{t}$ can be represented by $\mathbf{x}_{0}$:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}+f(\Delta\mathbf{h}_{t}).$ (10)
Through computation, $f(\mathbf{h}_{t})$ is defined as follows to meet the above two equations:
$f(\mathbf{h}_{t})=\sqrt{\bar{\alpha}_{t}}\mathbf{h}_{t}.$ (11)
Therefore, the forward process can be derived by incorporating the haze formation term into it for $T$ steps:

### v3

# Physics-guided Image Dehazing Diffusion

@@ Abstract
Due to the domain gap between real-world and synthetic hazy images, current data-driven dehazing algorithms trained on synthetic datasets perform well on synthetic data but struggle to generalize to real-world scenarios. To address this challenge, we propose Image Dehazing Diffusion Models (IDDM), a novel diffusion process that incorporates the atmospheric scattering model into noise diffusion. IDDM aims to use the gradual haze formation process to help the denoising Unet robustly learn the distribution of clear images from the conditional input hazy images. During the forward process, IDDM simultaneously introduces haze and noise into clear images, and then robustly separates them during the sampling process. By training with physics-guided information, IDDM shows the ability of domain generalization, and effectively restores the real-world hazy images despite being trained on synthetic datasets. Extensive experiments demonstrate the effectiveness of our method through both quantitative and qualitative comparisons with state-of-the-art approaches.
@@ 1 Introduction
Due to the attenuation of light in the atmosphere, images captured in hazy conditions often suffer from blurriness. These degradations not only affect visual quality but also significantly impair the performance of computer vision systems in critical applications such as autonomous driving and surveillance. To address the obvious harm to machine vision systems, extensive research [hazelines, dehazenet, nara2003, nara2000, NarasimhanNSK05, he_single_nodate] has been conducted in recent years on image dehazing.
Figure 1: Motivation for physics-guided Dehazing diffusion. Haze formation and the diffusion forward process share a common characteristic: both progressively degrade image clarity. IDDM integrates the physical haze formation process (depth-indexed) into diffusion (time-indexed) by simultaneously adding atmospheric scattering and noise.
Prior-based methods [he_single_nodate, hazelines, fattal, tan, zhou, zhoulight] frequently lack robustness in scenarios where prior information is unavailable. In contrast, data-driven deep learning methods [c2pnet, dehazenet, fang2025guided, fan2024driving, fan2025depth, FD-GAN] have proven to be more effective in handling complex environmental conditions. Current deep learning approaches for image dehazing [Zhang_2024_CVPR, OKNET] are often trained on synthetic datasets [reside, haze4k] with paired hazy and haze-free images, as acquiring real-world paired hazy data is challenging. However, models trained exclusively on synthetic data often exhibit poor generalization capabilities when applied to real-world images, especially under complex and variable weather conditions. This domain gap significantly limits the practical utility of these algorithms in real-world applications. Recent works have attempted to address this limitation through domain adaptation [DAD], physical priors [psd], or contrastive learning [Zheng_2023_CVPR], but the fundamental challenge remains.
@@
To implement this progressive learning approach, we turn to diffusion models, which have demonstrated remarkable success in various image generation and restoration tasks. We find that diffusion models offer an ideal foundation for this setting for several compelling reasons. The denoising diffusion models, which also contain inherent progressive learning strategies through their step-by-step noise addition and removal processes, excel in data distribution modeling, which can address the cross-domain challenge between synthetic and real images by learning the natural distribution of clear images. Moreover, there exists a remarkable parallel between the forward process of DDPM [DDPM] and the physical process of haze formation in real-world scenarios. In a forward diffusion process, a clear image gradually becomes pure noise over time, while in the haze formation process, it gradually accumulates haze as depth increases. Both processes share a similar trajectory: the image progressively deteriorates, either into noisy or hazy. This natural correspondence provides a theoretical basis for integrating physical models into generative frameworks and makes diffusion models particularly suitable as a foundation for image dehazing.
@@
To overcome this limitation, we utilize their similarities to deduce a novel diffusion model (IDDM) specifically for image dehazing, which explicitly models the formation of haze through the diffusion process. We model haze as a timestep-dependent haze component $\mathbf{h}_{t}$ to integrate physical principles into the diffusion process. A unique bi-directional enhancement framework is created around it, enabling denoising and dehazing processes to mutually promote each other. $\mathbf{h}_{t}$ forms the foundation of our novel image dehazing diffusion model and serves as a conditional vector guiding the diffusion model’s training, enabling accurate noise prediction across varying haze densities. Simultaneously, intermediate images with varying levels of noise and haze from the diffusion process are utilized to train the model to estimate haze under diverse conditions at each timestep. Our method demonstrates consistent improvements across multiple no-reference quality metrics on real-world datasets such as RTTS and BEDDE, validating its generalization ability to real-world conditions.
Notably, our formulation offers a key advantage over previous physics-guided methods. Unlike conventional physics-guided methods [Varghese2023] that separately estimate atmospheric light and transmission—leading to ill-posed multi-variable problems—our approach directly predicts the unified haze component $\mathbf{h}_{T}$ and implicitly recovers the attenuated part, circumventing contradictory estimations.
@@
(1) We introduce Image Dehazing Diffusion Models (IDDM), a novel framework that integrates the atmospheric scattering model into diffusion processes. This integration allows IDDM to simultaneously leverage the distribution learning capabilities of diffusion models and the physical constraints of atmospheric scattering, creating a physically-informed approach to dehazing.
(2) We propose a bidirectional training framework that mutually enhances diffusion model and haze estimator. The predicted haze component serves as conditioning to guide denosing process, while intermediate states generated by diffusion with varying noise-haze combinations provide step-wise supervision for training haze estimator.
@@ 2.1 Single Image Dehazing
Traditional image dehazing methods rely on hand-crafted priors such as dark channel [he_single_nodate] and haze-lines [hazelines], which often struggle with complex scattering scenarios. The emergence of deep learning shifted this paradigm, with early CNN-based approaches [dehazenet, ren2016, Zhang_2018_CVPR] learning to estimate transmission maps and atmospheric light, while later end-to-end methods [Dong_2020_CVPR, Li_2017_ICCV, Liu_2019_ICCV, Qin_Wang_Bai_Xie_Jia_2020, Qu_2019_CVPR] directly mapped hazy to clear images. However, models trained purely on data-driven mappings without physical constraints often produce artifacts and exhibit poor generalization to real-world scenarios.
To address these limitations, recent physics-guided approaches [Zheng_2023_CVPR, Wu_2023_CVPR, psd] integrate the atmospheric scattering model (ASM) to provide physical constraints during training. While these methods show improved performance, they typically require explicit estimation of multiple interdependent parameters (atmospheric light and transmission), leading to an ill-posed optimization problem. For real-world deployment, domain adaptation techniques [DAD, D4, ridcp] have been proposed to bridge the synthetic-to-real gap, though challenges remain in handling diverse weather conditions and dense haze scenarios. These findings motivate our investigation of diffusion models as a unified framework that naturally captures both the physical haze formation process and the distribution of real-world images.
@@ 2.2 Diffusion Models
Diffusion probabilistic models [pmlr-v37-sohl-dickstein15, DDPM] have emerged as powerful generative frameworks through their iterative denoising process, achieving remarkable success across diverse vision tasks including image restoration [Xie_2024_CVPR, 10021824, Garber_2024_CVPR], super-resolution [Gao_2023_CVPR], deblurring [Whang_2022_CVPR], and content generation [Rombach_2022_CVPR, Yang_2023_CVPR, Zhang_2023_CVPR]. Score-based formulations [NEURIPS2019_3001ef25, meng2022sdedit, Chung_2022_CVPR] provide an alternative perspective, enabling flexible conditional generation through various guidance mechanisms.
Despite their success in general image processing, applying diffusion models to physically-constrained restoration tasks like dehazing remains underexplored. Recent unified multimodal models [blip3o] demonstrate the effectiveness of bi-directional conditioning between different model components, where intermediate representations mutually supervise each task. Inspired by these advances, we propose integrating physical haze formation into the diffusion framework, where the progressive nature of diffusion naturally aligns with atmospheric scattering processes.
Figure 2: Overview of the proposed framework. We introduce a physics-guided diffusion process that explicitly models haze formation through the atmospheric scattering model. (A) During forward diffusion, both physical haze $\mathbf{h}_{t}$ and noise $\boldsymbol{\epsilon}_{t}$} are progressively added to clear images through shared timesteps, where $\mathbf{h}_{t}$ accumulates according to depth-dependent scattering attenuation. (B) The reverse process leverages a Haze Estimator (HtNet) to predict timestep-dependent haze intensity, which conditions the denoising network for physics-aware image restoration. (C) A bi-directional training strategy is employed: the frozen denoising U-Net supervises HtNet training via consistency regularization, while predicted haze components guide noise prediction across varying degradation levels.
@@ 3 Image Dehazing Diffusion Models
3.1 Why Diffusion for Dehazing
Image dehazing aims to recover a clear image $J$ from its hazy observation $I$. According to the atmospheric scattering model (ASM) [nara2000, nara2003], this degradation process is described as:
$I$ $=\underbrace{Je^{-\sigma z}}_{\text{Attenuation}}+\underbrace{\int_{0}^{z}Ae^{-\sigma z}dz}_{\text{Haze Component}}$ (1)
$=\underbrace{JT_{r}}_{\text{Attenuation}}+\underbrace{A^{\prime}(1-T_{r})}_{\text{Haze Component}}.$
where $A=\frac{A^{\prime}}{\sigma},T_{r}=e^{-\sigma z}$. Here, $J$ denotes a clear image, $I$ denotes a hazy image, $A$ denotes the airlight, $z$ denotes depth, and $T_{r}$ denotes the transmission map. Here, a hazy image can be divided into two parts: attenuation and haze component.
Base on Equation 1, the recovery of the clear image $J$ necessitates two steps: calculating the Haze Component and inferring the transmission map $T_{r}$. A key observation is that the haze component accumulates progressively with depth – as $z$ increases from 0 to infinity, the image gradually transitions from clear $J$ to completely hazy $A^{\prime}$.
This progressive nature of haze formation shares a remarkable similarity with diffusion models, which also involve gradual transformation over time steps. In diffusion models, a clear image progressively becomes noise through iterative addition. Given a data point sampled from a real data distribution $\mathbf{x}_{0}$, the forward diffusion process of DDPM at step $t$ can be explained as:
@@ 3.1 Why Diffusion for Dehazing
where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\alpha_{i}$ denotes signal weight. In this equation, $\mathbf{x}_{t}$ represents the data at time step $t$ of diffusion, $\alpha_{t}$ is the scaling factor that controls noise, and $\boldsymbol{\epsilon}_{t}$ is Gaussian noise sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. Both processes share a fundamental similarity: progressive degradation through additive accumulation. While diffusion models add Gaussian noise $\boldsymbol{\epsilon}_{t}$ at each timestep, haze formation adds atmospheric scattering at each depth increment.
We ask: Can we leverage this similarity to design a physics-guided diffusion model that naturally incorporates the haze formation process? Recent works [domain1, zhang2023domain] have demonstrated the effectiveness of diffusion models for domain adaptation tasks, suggesting their inherent capability to learn transferable representations across different domains. This provides additional theoretical motivation for applying diffusion models to address the synthetic-to-real generalization challenge in image dehazing. Traditional dehazing methods either estimate $T_{r}$ and A separately or learn direct mappings, but they struggle with domain generalization. By integrating the physical process into the diffusion framework, we aim to learn not just the mapping but the underlying formation process, leading to better generalization.
3.2 Modeling Haze Formation as a Diffusion Process
Motivated by the analogy between progressive haze accumulation and diffusion noise addition, we reformulate haze formation as a time-indexed process that mirrors diffusion timesteps.
From Depth-Dependent to Time-Dependent. In the atmospheric scattering model, the amount of haze observed at a pixel $\mathbf{h}(x,y)$ is determined by integrating atmospheric scattering along the light path from depth 0 to the scene depth $Z(x,y)$:
$\mathbf{h}(x,y)=A\int_{0}^{Z(x,y)}e^{-\sigma z}dz.$ (3)
We reformulate this continuous, depth-dependent process as a discrete, time-dependent process. For any pixel with scene depth $Z$, we can partition the total integration path $[0,Z]$ into $T$ sequential segments. The total haze $\mathbf{h}(Z)$ can then be expressed as the sum of the haze contributed by each segment $t$, where each segment has a depth increment of $\frac{Z}{T}$:
$\mathbf{h}(Z)=A\sum_{t=1}^{T}\int_{\frac{t-1}{T}\cdot Z}^{\frac{t}{T}\cdot Z}e^{-\sigma z}dz.$ (4)
Since the sum of integrals equals the integral of the sum, this discrete formulation in Equation (4) is mathematically equivalent to the continuous formulation in Equation (3).
Per-Timestep Haze Integration. This $T$-step reformulation allows us to define the progressive state of haze accumulation. We define $\mathbf{h}_{t}$ as the total haze accumulated up to the $t$-th timestep, corresponding to the integration from depth 0 to $\frac{t}{T}\cdot Z$:
$\mathbf{h}_{t}=\int_{0}^{\frac{t}{T}\cdot Z}Ae^{-\sigma z}dz.$ (5)
The incremental haze added between consecutive timesteps t1t-1 and $t$ (i.e., from depth $\frac{t-1}{T}\cdot Z$ to $\frac{t}{T}\cdot Z$) is:
$\Delta\mathbf{h}_{t}=\mathbf{h}_{t}-\mathbf{h}_{t-1}=A\int_{\frac{t-1}{T}\cdot Z}^{\frac{t}{T}\cdot Z}e^{-\sigma z}dz.$ (6)
Consequently, the final haze $\mathbf{h}_{T}$ at depth $Z$ is the cumulative sum of all incremental haze components:
$\mathbf{h}_{T}=\sum_{t=1}^{T}\Delta\mathbf{h}_{t}.$ (7)
The parallel with diffusion. The above formulation reveals that haze accumulation is inherently additive, similar to how noise accumulates in diffusion models. In the end, we transfer haze from a depth-dependent value into a time-dependent value of diffusion. Therefore, both processes progressively degrade image quality through the increase of time step; both involve accumulation over time, making them naturally compatible for unified modeling in a single framework.
3.3 The IDDM Forward Process
Based on Equation 1, we reformulate the diffusion forward process to adhere to physical principles: we set the base state $\mathbf{x}_{0}$ as the attenuated clear image $JT_{r}$, and restore attenuation $T_{r}$ later based on $\mathbf{h}_{t}$. The diffusion process then progressively adds the haze component $\mathbf{h}_{T}$ to $\mathbf{x}_{0}$, directly modeling the physical accumulation process $I=\mathbf{x}_{0}+\mathbf{h}_{T}$.
Our key design simultaneously conducts haze diffusion and noise diffusion through shared timesteps (Fig. 2A). At each step $t$, haze component $\mathbf{h}_{t}$ accumulates according to the atmospheric scattering model (Eq. 5), while Gaussian noise $\boldsymbol{\epsilon}_{t}$ is added following the standard diffusion schedule. This synchronization enables the model to learn the joint distribution of haze and noise at varying degradation levels.
However, while both diffusion and haze accumulation exhibit progressive accumulation over time, they follow different mathematical formulations. Standard diffusion applies a linear combination with coefficient $\sqrt{\bar{\alpha}_{t}}$, whereas haze accumulation accumulates $\Delta\mathbf{h}_{t}$ without it. Therefore, integrating haze into the diffusion framework requires careful mathematical formulation to ensure compatibility. To maintain mathematical consistency, we derive the appropriate scaling for incremental haze addition (detailed derivation in Appendix), and propose a modified forward process that incorporates haze in a diffusion-compatible manner: